# FedMentor: Domain-Aware Differential Privacy for Heterogeneous Federated LLMs in Mental Health

### v1

@@ 5 Main results
Privacy and safety. FedMentor enforces strict per-domain DP while achieving safer outputs and nearly the same utility as non-private FL (Table 1). Across Dreaddit, IRF, and MultiWD, TSR increases and toxicity decreases, while BERTScore F1 and ROUGE-L shift only slightly. On MultiWD (few-shot, ParetoQ-350M), TSR improves from 93.0% to 95.0% and mean toxicity drops from 2.92 to 1.98. On IRF, the strictest privacy domain, FedMentor achieves 92% safe outputs versus 96% for FL and keeps BERTScore F1 and ROUGE-L within 0.1 to 0.2 of the no-DP baseline. Scaling up, FedMentor with Qwen3-0.6B matches the FL safe rate ($\approx$93%) with identical BERTScore F1, and with Qwen3-1.7B the remaining utility gap narrows further. Overall, FedMentor provides per-domain private training that improves safety over vanilla FL and approaches the centralized upper bound; Appendix C reports consistent trends for SmolLM2-360M/1.7B.
Figure 3: ParetoQ-350M: (a) REL and (b) B-F1 with TSR under Baseline, Static, and Uniform. Qwen3-1.7B: (c) $\varepsilon$ ablation of B-F1, TSR, and REL on global evaluation and on Dreaddit, IRF, and MultiWD. All panels report the final global model after 8 rounds (2 local epochs per round).
Utility under heterogeneity. FedMentor sustains competitive utility on non-IID data, and few-shot fine-tuning consistently outperforms zero-shot prompting. In Table 1, under few-shot training FedMentor’s performance closely matches FL without DP: for example, on Dreaddit the FedMentor BERTScore F1 and ROUGE-L are within 0.1 and 0.2 of the no-DP FL values, and safe rate remains high (98% vs 99%). Few-shot adaptation yields clear gains over zero-shot across all models. In particular, relevance and BERTScore improve by roughly 1–3 points with few-shot training compared to zero-shot (e.g., IRF relevance rises from 69.8% to 71.4% under FedMentor with Qwen3-0.6B). We observe no domain collapse: Figure 2 shows only modest variation across datasets. Furthermore, at the client level FedMentor exhibits consistent behavior with FL: Figure 6 (Appendix C) shows per-client relevance scores overlapping (within $\pm$1%), indicating uniform utility across clients. Although absolute metrics are lower without fine-tuning, the relative gap between FedMentor and FL remains stable, reinforcing that strong privacy comes at minimal utility cost under heterogeneous data.
@@
Baseline Domain Spec. 2.0 / 0.5 / 1.5 derived $97.33$ $94$ $100$ $2.49$ 66
IRF Sensitivity Domain Spec. fixed 2.0 / 0.1 / 1.5 derived $78.33$ $66$ $88$ $9.18$ $22$
@@
Utility Threshold Global B-F1 $\tau{=}0$ $74.67$ $65$ $85$ $8.18$ 2020
Round Client 0 (GB) Client 1 (GB) Client 2 (GB)
0 $18.85$ $20.04$ $21.23$
1 $22.42$ $23.60$ $24.79$
2 $25.98$ $27.16$ $28.35$
3 $29.54$ $30.72$ $31.91$
4 $33.10$ $34.29$ $35.47$
5 $36.66$ $37.85$ $39.03$
6 $40.22$ $41.41$ $42.59$
7 $43.78$ $44.97$ $46.16$
Practical efficiency. FedMentor satisfies practical efficiency requirements in communication, memory, and speed on single-GPU clients. Table 2 reports compact LoRA updates: 16.56 MB for ParetoQ-350M, 38.50 MB for Qwen3-0.6B, and 66.50 MB for Qwen3-1.7B. With three clients, this translates to only about 49.7 MB, 110.0 MB, and 172.9 MB communicated per round, since only adapters (not full model weights) are exchanged. Peak memory per client remains about 33.7 GB for the 350M model and 77.9 GB for the Qwen3 models, within a single 80 GB GPU; per-round Qwen3-1.7B memory for the IRF $\varepsilon$ sweep in Table 4 shows all clients below the device limit. Average round time (two local epochs, three clients) is 7.6, 10.8, and 11.4 minutes for the 350M, 0.6B, and 1.7B backbones. Dataset trends follow size: IRF is fastest (3.96 min at 350M), Dreaddit is intermediate, and MultiWD is slowest (up to 12.3 min at 1.7B). Communication cost is constant across datasets (50–173 MB) as it depends on model size only. Together with Table 1, these results indicate low overhead and practical training times.
@@ 6 Ablation studies
Uniform privacy. Uniform Privacy applies a single global privacy budget to all domains. In Figure 3a–b and Table 3, this choice markedly improves fairness in safety: with a uniform $\epsilon{=}1.0$, the Toxicity Safe Rate (TSR) concentrates at $\sim$99–100% and the client spread contracts to 1% (vs. 6% under the domain-specific baseline). In contrast, the Static Noise variant shows lower average safety (87.33% TSR) and a larger spread (18%). The utility-gated variant with $\tau{=}0$ yields the lowest safety (74.67% TSR). BERTScore F1 and REL remain close to the baseline across settings, indicating that the fairness gains from a uniform budget come with only modest utility change. All panels in Figure 3 report the final global model after 8 rounds.
@@
See Appendix D.1 for the utility-threshold ablation and Appendix D.2 for per-client memory scaling across rounds.
7 Related work
LLMs in mental health. LLMs are rapidly adopted for mental health applications, supporting tasks such as condition detection, diagnosis, and therapeutic dialogue Wang et al. (2021); Yang et al. (2023b, a); Xu et al. (2024); Wang et al. (2024b); Zheng et al. (2025); Chan et al. (2025); Hengle et al. (2024); Haider et al. (2025). While these systems demonstrate strong potential for supportive interaction, most current approaches are centralized and depend on aggregating sensitive dialogue data. Such data are scarce due to confidentiality concerns, and strict regulations, including HIPAA and GDPR, further constrain sharing Kwesi et al. (2025); Baidal et al. (2025); Nguyen et al. (2025). As a result, datasets are often small and biased, which limits the robustness and generalization of models.
These limitations reveal a critical privacy–utility trade-off that existing LLM methods rarely address, reducing their feasibility in real-world clinical contexts Lin et al. (2023); Chung et al. (2023); Gabriel et al. (2024); Kumar et al. (2024). FedMentor addresses this challenge by integrating Federated Learning with domain-aware Differential Privacy, enabling collaborative training without centralizing data. This framework allows models to learn from sensitive mental health texts while preserving confidentiality, thereby enhancing both trust and clinical applicability.
@@ 7 Related work
FL for mental health. FL has demonstrated performance comparable to centralized training while satisfying strict privacy regulations such as HIPAA and GDPR Lieftink et al. (2024); Peng et al. (2024); Woisetschläger et al. (2024). By training directly on decentralized data, FL is well-suited for mobile health and multi-clinic settings. In mental health, FedTherapist Shin et al. (2023) leverages on-device FL with speech and keyboard signals to track depression, stress, and mood, and FedMood Xu et al. (2021) introduces a multi-view framework for depression diagnosis using heterogeneous mobile health data. These studies, along with work on mobile sensing, electronic health records, and multi-institutional collaboration, demonstrate the potential of FL for sensitive mental health applications Rauniyar et al. (2023).
Building on these foundations, recent work has extended FL to LLMs for conversational and diagnostic support. FedMentalCare Sarwar (2025) combines FL with LoRA to fine tune lightweight LLMs while reducing communication costs. However, many methods assume homogeneous data and only partly address scalability and efficiency. We introduce FedMentor, which enables heterogeneous FL through explicit domain modeling, LoRA-based communication reduction, and domain-aware Differential Privacy for practical and privacy-preserving LLM fine tuning in mental health.
Extended discussion of Differential Privacy in FL appears in App. A.
@@ 8 Conclusion
We presented FedMentor, a federated fine-tuning framework that combines domain-aware Differential Privacy with LoRA adapters for sensitive mental health use. Across heterogeneous domains, FedMentor improves safety while keeping utility close to non-private Federated Learning and near centralized training. Practicality follows from communicating adapter weights only, which keeps memory and bandwidth within single GPU budgets. Ablation studies show that domain-specific budgets and adaptive noise control are essential, since removing either lowers accuracy and widens client-level disparities. This study has limits: a restricted set of datasets, no explicit audit of demographic or linguistic bias, and reliance on automatic metrics without clinician review. Future work will scale to larger and more diverse federations, incorporate clinician-in-the-loop and multilingual evaluation, and extend fairness analyses to identify and mitigate potential biases.
Appendix A Additional related work
Differential privacy in FL. Differential privacy (DP) has become a standard approach for protecting sensitive data in ML, including FL Dwork et al. (2014); Abadi et al. (2016); Girgis et al. (2021); Noble et al. (2022). Applying DP in practice, however, remains difficult, particularly in mental health settings. Traditional methods that uniformly clip gradients and add noise to updates often degrade performance in healthcare, where datasets are small and signals are fine-grained McMahan et al. (2018). Although DP has been studied in areas such as electronic health records and medical imaging, its application in mental health remains limited, with only a few recent studies exploring privacy-by-design strategies Choudhury et al. (2019); Sawhney et al. (2023); Romijnders et al. (2024). Recent research further highlights that not all model components or data are equally sensitive, motivating personalized and domain-aware DP strategies that adapt privacy to data characteristics Ebadi et al. (2015); Jorgensen et al. (2015); Liu et al. (2021, 2022, 2024b). In LLMs, parameter-efficient methods such as DP-LoRA Liu et al. (2025) reduce risk and communication costs by perturbing only low-rank adapter updates. However, most approaches still adopt a one-size-fits-all noise and overlook the unique sensitivities of mental health text. To overcome this limitation, we introduce FedMentor, a framework that integrates domain-aware DP with heterogeneous federated learning to provide strong privacy guarantees while preserving clinical utility.
@@ B.2 Model specification
MobileLLM-ParetoQ (350M) Liu et al. (2024a). MobileLLM targets on device efficiency with a deep thin Transformer, embedding tying, grouped query attention, and blockwise weight sharing. ParetoQ adds sub 4 bit quantization with quantization aware training and learnable scaling. In FedMentor, we load 4 bit NF4 with bf16 compute and fine-tune only LoRA adapters (rank 8), producing very small adapter states. Domain aware DP is applied directly to adapters, which keeps privacy cost and per round communication low.
@@ B.3 Baseline specification
Dataset-specific instruction prompts Dreaddit. Analyze this Reddit post for signs of psychological stress. Provide a brief, supportive response in 1–2 sentences. If stress is present, acknowledge feelings and suggest one helpful action. If no stress, provide encouragement. Be empathetic but avoid medical advice.
Post: {post}
IRF. Analyze this post for signs of interpersonal risk (thwarted belongingness or perceived burdensomeness). Provide a brief, supportive response in 1–2 sentences. Be empathetic but avoid medical advice.
Post: {post}
MultiWD. Analyze this post for social and emotional wellness concerns. Social wellness involves relationships, connection, and community. Emotional wellness involves managing feelings, stress, and mental health. Provide a brief, supportive response in 1–2 sentences. Be empathetic and suggest practical coping strategies.
Post: {post}
@@
(a) Zero-shot model wrappers Qwen
@@
(b) Few-shot and train-time shape Few-shot wrapper (generic $k$).
@@
Figure 5: Prompt templates: (a) Zero-shot wrappers for Qwen, SmolLM2, and MobileLLM; (b) Few-shot wrapper and exact train-time prompt shape.
@@ B.4 Detailed evaluation metrics
Centralized (w/o FL). We evaluate toxicity mean and toxicity max of generated responses using Detoxify, safe rate (fraction of responses below a fixed toxicity threshold), BERTScore F1 for semantic similarity, relevance mean via embedding based cosine similarity, ROUGE-L for lexical overlap, and perplexity mean as a fluency proxy. These quantify response quality and safety for single model generation.
Federated (FL). For each client and in aggregate we report toxicity mean, safe rate, BERTScore F1, relevance mean, and ROUGE-L. In accordance with the code, toxicity max and perplexity mean are not computed in the federated evaluator.
FL w/ DP (system metrics). To characterize privacy–utility tradeoffs and overhead, we log domain sensitivity, training time, final train loss, final eval loss, memory used (MB), and LoRA weights size (MB) per client, along with per round aggregates: average train loss, average eval loss, total training time, average memory usage, and communication overhead (MB). Communication overhead reflects the size of adapter payloads, since only LoRA adapters are shared rather than full model weights.
B.5 Prompt construction for zero-shot and few-shot
Zero-shot prompting. We follow standard zero-shot evaluation, supplying a single instruction plus the post, wrapped in the backbone native template: Qwen uses chat role tags, SmolLM2 uses Input/Response headers, and MobileLLM uses plain text with a trailing Response:. Instructions are domain aligned: stress identification and supportive coping for Dreaddit; empathetic interpersonal risk screening for IRF (thwarted belongingness or perceived burdensomeness); and brief practical guidance for social and emotional wellness cues in MultiWD. This uniform yet domain aware setup isolates instruction following while keeping comparisons consistent across datasets and models. Exact wrappers and examples are shown in Figure 5(a), with dataset specific instructions in Figure 4.
Few-shot prompting. We implement few-shot as supervised LoRA adaptation on each client rather than in context exemplars. Only adapters are trained on a quantized backbone; updates are privatized with domain aware Gaussian noise and aggregated with FedAvg. Short supportive targets align tone and structure to each domain, which increases safe rate, reduces toxicity mean, and improves BERTScore F1. At inference, we reuse the zero-shot wrapper for a clean comparison. The generic train time wrapper and target concatenation appear in Figure 5(b).
@@ B.5 Prompt construction for zero-shot and few-shot
B.6 Hardware details.

### v2

@@ 5 Main results
Privacy and Safety. FedMentor enforces strict per-domain DP while achieving safer outputs and nearly the same utility as non-private FL (Table 1). Across Dreaddit, IRF, and MultiWD, TSR increases and toxicity decreases, while BERTScore F1 and ROUGE-L change slightly. On MultiWD (few-shot, ParetoQ-350M), TSR improves from 93.0% to 95.0% and mean toxicity drops from 2.92 to 1.98. On IRF, the strictest privacy domain, FedMentor achieves 92% safe outputs versus 96% for FL and keeps BERTScore F1 and ROUGE-L within 0.1 to 0.2 of the no-DP baseline. Scaling up, FedMentor with Qwen3-0.6B matches the FL safe rate ($\approx$93%) with identical BERTScore F1, and with Qwen3-1.7B the utility gap narrows further. Overall, FedMentor provides per-domain private training that improves safety over vanilla FL and approaches the centralized upper bound; App. C reports consistent trends for SmolLM2-360M/1.7B.
Figure 4: ParetoQ-350M: (a) REL and (b) B-F1 with TSR under Baseline, Static, and Uniform. Qwen3-1.7B: (c) $\varepsilon$ ablation of B-F1, TSR, and REL on global evaluation and on Dreaddit, IRF, and MultiWD. All panels report the final global model after 8 rounds (2 local epochs per-round).
Utility under Heterogeneity. FedMentor sustains competitive utility on non-IID data, and few-shot fine-tuning consistently outperforms zero-shot prompting. In Table 1, under few-shot training FedMentor’s performance closely matches FL without DP: for example, on Dreaddit the FedMentor BERTScore F1 and ROUGE-L are within 0.1 and 0.2 of the no-DP FL values, and safe rate remains high (98% vs 99%). Few-shot adaptation yields clear gains over zero-shot across all models. In particular, relevance and BERTScore improve by roughly 1–3 points with few-shot training compared to zero-shot (e.g., IRF relevance rises from 69.8% to 71.4% under FedMentor with Qwen3-0.6B). We observe no domain collapse: Figure 3 shows only modest variation across datasets. Furthermore, at the client level FedMentor exhibits consistent behavior with FL: Figure 5 (App. C) shows per-client relevance scores overlapping (within $\pm$1%), indicating uniform utility across clients. Although absolute metrics are lower without fine-tuning, the relative gap between FedMentor and FL remains stable, reinforcing that strong privacy comes at minimal utility cost under heterogeneous data.
@@
Baseline
Domain
Spec.
2.0 / 0.5 / 1.5 derived $97.33$ $94$ $100$ $2.49$ 66
IRF Sensitivity
Domain
Spec.
fixed 2.0 / 0.1 / 1.5 derived $78.33$ $66$ $88$ $9.18$ $22$
@@
Utility Threshold Global
B-F1
$\tau{=}0$
$74.67$ $65$ $85$ $8.18$ 2020
Practical Efficiency. FedMentor satisfies practical efficiency requirements in communication, memory, and speed on single-GPU clients. Table 2 reports compact LoRA updates: 16.56 MB for ParetoQ-350M, 38.50 MB for Qwen3-0.6B, and 66.50 MB for Qwen3-1.7B. With three clients, this translates to about 49.7 MB, 110.0 MB, and 172.9 MB communicated per-round, since only adapters, not full weights, are exchanged. Peak memory per client remains about 33.7 GB for the 350M model and 77.9 GB for the Qwen3 models, within a single 80 GB GPU; per-round Qwen3-1.7B memory for the IRF $\varepsilon$ sweep in Table 5 (App. C) shows all clients below the device limit. Average round time (two local epochs, three clients) is 7.6, 10.8, and 11.4 minutes for the 350M, 0.6B, and 1.7B backbones. Dataset trends follow size: IRF is fastest (3.96 min at 350M), Dreaddit is intermediate, and MultiWD is slowest (up to 12.3 min at 1.7B). Communication cost is constant across datasets (50–173 MB) as it depends on model size only. Together with Table 1, these results indicate low overhead and practical training times.
@@ 6 Ablation Studies
Uniform Privacy. Uniform Privacy applies a single global privacy budget to all domains. In Figure 4a–b and Table 3, this choice markedly improves fairness in safety: with a uniform $\epsilon{=}1.0$, the TSR concentrates at $\sim$99–100% and the client spread contracts to 1% (vs. 6% under the domain-specific baseline). In contrast, the static noise variant shows lower average safety (87.33% TSR) and a larger spread (18%). The utility-gated variant with $\tau{=}0$ yields the lowest safety (74.67% TSR). BERTScore F1 and REL remain close to the baseline across settings, indicating that the fairness gains from a uniform budget come with only modest utility change. All panels in Figure 4 report the final global model after 8 rounds.
@@
Utility Threshold Calibration ($\tau$). We evaluate a utility-driven noise adjustment where the server reduces noise only if B-F1 falls below a threshold $\tau$. Table 3 shows that the $\tau{=}0$ setting yields the lowest safety, with TSR at 74.67% and a spread of 20%, compared to the baseline TSR of 97.33% and spread of 6%. The observation is consistent with the ParetoQ-350M trends in Figure 4a–b: aggressive utility gating undermines safety, indicating the need for a calibrated threshold.
Per-Client Memory Scaling across Rounds Table 5 reports memory usage per client for Qwen3-1.7B under the IRF $\epsilon{=}1.0$ setting. Memory increases smoothly over rounds: at round 0, clients use about 18.85, 20.04, and 21.23 GB; by round 7, these reach 43.78, 44.97, and 46.16 GB. The monotonic rise indicates predictable scaling during training while keeping per-client usage within a narrow band across clients.
7 Related Works
Round Client 0 (GB) Client 1 (GB) Client 2 (GB)
0 $18.85$ $20.04$ $21.23$
1 $22.42$ $23.60$ $24.79$
2 $25.98$ $27.16$ $28.35$
3 $29.54$ $30.72$ $31.91$
4 $33.10$ $34.29$ $35.47$
5 $36.66$ $37.85$ $39.03$
6 $40.22$ $41.41$ $42.59$
7 $43.78$ $44.97$ $46.16$
LLMs in Mental Health. LLMs are rapidly adopted for mental health applications, supporting tasks such as condition detection, diagnosis, and therapeutic dialogue Wang et al. (2021); Yang et al. (2023b, a); Xu et al. (2024); Wang et al. (2024a); Mohammadi et al. (2024); Zheng et al. (2025); Chan et al. (2025); Hengle et al. (2024); Haider et al. (2025). While these systems demonstrate strong potential for supportive interaction, most current approaches are centralized and depend on aggregating sensitive dialogue data. Such data are scarce due to confidentiality concerns, and strict regulations, including HIPAA and GDPR, further constrain sharing Kwesi et al. (2025); Baidal et al. (2025); Nguyen et al. (2025). As a result, datasets are often small and biased, which limits the robustness and generalization of models.
These limitations reveal a critical privacy-utility trade-off that existing LLM methods rarely address, reducing their feasibility in real-world clinical contexts Lin et al. (2023); Chung et al. (2023); Gabriel et al. (2024); Kumar et al. (2024). FedMentor addresses this challenge by integrating Federated Learning with domain-aware Differential Privacy, enabling collaborative training without centralizing data. This framework allows models to learn from sensitive mental health texts while preserving confidentiality, thereby enhancing both trust and clinical applicability.
FL for Mental Health. Federated Learning (FL) has demonstrated performance comparable to centralized training while satisfying strict privacy regulations such as HIPAA and GDPR Lieftink et al. (2024); Peng et al. (2024); Woisetschläger et al. (2024). By training directly on decentralized data, FL is well-suited for mobile health and multi-clinic settings. In mental health, FedTherapist Shin et al. (2023) leverages on-device FL with speech and keyboard signals to track depression, stress, and mood, and FedMood Xu et al. (2021) introduces a multi-view framework for depression diagnosis using heterogeneous mobile health data. These studies, along with work on mobile sensing, electronic health records, and multi-institutional collaboration, demonstrate the potential of FL for sensitive mental health applications Rauniyar et al. (2023).
Building on these foundations, recent work has extended FL to LLMs for conversational and diagnostic support. FedMentalCare Sarwar (2025) combines FL with LoRA to fine tune lightweight LLMs while reducing communication costs. However, many methods assume homogeneous data and only partly address scalability and efficiency. We introduce FedMentor, which enables heterogeneous FL through explicit domain modeling, LoRA-based communication reduction, and domain-aware DP for practical and privacy-preserving LLM fine tuning in mental health.
App. A discusses related work on Differential Privacy in FL.
@@ 8 Conclusion
We introduced FedMentor, a federated fine-tuning framework that pairs domain-aware Differential Privacy with LoRA adapters for sensitive mental health dialogue. Across heterogeneous domains, FedMentor raises safety while keeping utility near non-private Federated Learning and near centralized training. Practicality follows from communicating adapter weights only, which keeps memory and bandwidth within single GPU budgets. Ablations confirm that domain-specific budgets and adaptive noise are essential, since removing either reduces accuracy and amplifies client disparities. The study utilizes a limited set of datasets and automated metrics without clinician review. Future work will scale federations, incorporate clinician-in-the-loop approaches, and conduct multilingual evaluation, as well as extend fairness analyses.
Limitations
Several factors limit the present study. First, the evaluation utilizes three English social-media datasets, treating each dataset as a single client, which limits external validity for larger cross-silo federations and clinical conversations. Second, assessment relies on automatic quality and safety proxies, including detector-based toxicity and embedding similarity, without clinician review or user studies. Third, the work reports no audit of demographic or linguistic subgroups and does not provide round-level composition accounting or attack-based testing for privacy loss. Fourth, system results reflect an A100 80 GB environment and a small number of clients, so hardware and scale shifts may alter efficiency. Finally, comparisons do not include alternative differentially private fine-tuning strategies beyond LoRA-based adapters.
Appendix A Additional Related Works
Differential Privacy in FL. Differential privacy (DP) has become a standard approach for protecting sensitive data in ML, including FL Dwork et al. (2014); Abadi et al. (2016); Girgis et al. (2021); Noble et al. (2022). Applying DP in practice, however, remains difficult, particularly in mental health settings. Traditional methods that uniformly clip gradients and add noise to updates often degrade performance in healthcare, where datasets are small and signals are fine-grained McMahan et al. (2018). Although DP has been studied in areas such as electronic health records and medical imaging, its application in mental health remains limited, with only a few recent studies exploring privacy-by-design strategies Choudhury et al. (2019); Sawhney et al. (2023); Romijnders et al. (2024). Recent research further highlights that not all model components or data are equally sensitive, motivating personalized and domain-aware DP strategies that adapt privacy to data characteristics Ebadi et al. (2015); Jorgensen et al. (2015); Liu et al. (2021, 2022, 2024a). In LLMs, parameter-efficient methods such as DP-LoRA Liu et al. (2025) reduce risk and communication costs by perturbing only low-rank adapter updates. However, most approaches still adopt a one-size-fits-all noise and overlook the unique sensitivities of mental health text. To overcome this limitation, we introduce FedMentor, a framework that integrates domain-aware DP with heterogeneous federated learning to provide strong privacy guarantees while preserving clinical utility.
@@ B.2 Model Specification
MobileLLM-ParetoQ (350M) Liu et al. (2024b). MobileLLM targets on device efficiency with a deep thin Transformer, embedding tying, grouped query attention, and blockwise weight sharing. ParetoQ adds sub 4 bit quantization with quantization aware training and learnable scaling. In FedMentor, we load 4 bit NF4 with bf16 compute and fine-tune only LoRA adapters (rank 8), producing very small adapter states. Domain aware DP is applied directly to adapters, which keeps privacy cost and per round communication low.
@@ B.3 Baseline Specification
Dataset-Specific Instruction Prompts Dreaddit. Analyze this Reddit post for signs of psychological stress. Provide a brief, supportive response in 1–2 sentences. If stress is present, acknowledge feelings and suggest one helpful action. If no stress, provide encouragement. Be empathetic but avoid medical advice. Post: {post} IRF. Analyze this post for signs of interpersonal risk (thwarted belongingness or perceived burdensomeness). Provide a brief, supportive response in 1–2 sentences. Be empathetic but avoid medical advice. Post: {post} MultiWD. Analyze this post for social and emotional wellness concerns. Social wellness involves relationships, connection, and community. Emotional wellness involves managing feelings, stress, and mental health. Provide a brief, supportive response in 1–2 sentences. Be empathetic and suggest practical coping strategies. Post: {post}
@@
Zero-Shot Model Wrappers Qwen
@@
Figure 7: Prompt templates: Zero-shot wrappers for Qwen, SmolLM2, and MobileLLM.
Few-Shot and Train-Time Shape Few-shot wrapper (generic $k$).
@@
Figure 8: Prompt templates: Few-shot wrapper and exact train-time prompt shape.
@@ B.4 Detailed Evaluation Metrics
Centralized (w/o FL, w/o DP). We evaluate toxicity mean and toxicity max of generated responses using Detoxify, safe rate (fraction of responses below a fixed toxicity threshold), BERTScore F1 for semantic similarity, relevance mean via embedding based cosine similarity, ROUGE-L for lexical overlap, and perplexity mean as a fluency proxy. These quantify response quality and safety for single model generation.
Federated with LoRA (w/o DP). For each client and in aggregate we report toxicity mean, safe rate, BERTScore F1, relevance mean, and ROUGE-L. In accordance with the code, toxicity max and perplexity mean are not computed in the federated evaluator.
Federated with LoRA (w/ DP) (System Metrics). To characterize privacy–utility tradeoffs and overhead, we log domain sensitivity, training time, final train loss, final eval loss, memory used (MB), and LoRA weights size (MB) per client, along with per round aggregates: average train loss, average eval loss, total training time, average memory usage, and communication overhead (MB). Communication overhead reflects the size of adapter payloads, since only LoRA adapters are shared rather than full model weights.
@@
B.5 Prompt Construction for Zero-Shot and Few-Shot
Zero-Shot Prompting. We follow standard zero-shot evaluation, supplying a single instruction plus the post, wrapped in the backbone native template: Qwen uses chat role tags, SmolLM2 uses Input/Response headers, and MobileLLM uses plain text with a trailing Response:. Instructions are domain aligned: stress identification and supportive coping for Dreaddit; empathetic interpersonal risk screening for IRF (thwarted belongingness or perceived burdensomeness); and brief practical guidance for social and emotional wellness cues in MultiWD. This uniform yet domain aware setup isolates instruction following while keeping comparisons consistent across datasets and models. Exact wrappers and examples are shown in Figure 7, with dataset specific instructions in Figure 6.
Few-Shot Prompting. We implement few-shot as supervised LoRA adaptation on each client rather than in context exemplars. Only adapters are trained on a quantized backbone; updates are privatized with domain aware Gaussian noise and aggregated with FedAvg. Short supportive targets align tone and structure to each domain, which increases safe rate, reduces toxicity mean, and improves BERTScore F1. At inference, we reuse the zero-shot wrapper for a clean comparison. The generic train time wrapper and target concatenation appear in Figure 8.
B.6 Hardware Details